# Space spanned by characteristic exponents

### v2

@@ 2.3. Transcendental points
Write $h=g\circ\psi_{\mathbb{C}}$, where $g\in\mathcal{O}^{L}(R(X))\otimes_{L}K$. Write $g=\pi_{L}^{*}(g_{1})+\pi_{L}^{*}(g_{2})i$, where $g_{1},g_{2}\in\mathcal{O}(R(X_{K}))$. Since $h(x)\in K$, we see that both $\pi_{L}^{*}(g_{1})(\psi_{\mathbb{C}}(x))$ and $\pi_{L}^{*}(g_{2})(\psi_{\mathbb{C}}(x))$ are in $L$. The map $\pi_{L}|_{\psi_{\mathbb{C}}(x)}:\psi_{\mathbb{C}}(x)\to Y(x)_{L}$ induces an embedding
$(\pi_{L}|_{\psi_{\mathbb{C}}(x)})^{*}:\mathcal{O}(Y(x)_{L})\hookrightarrow\mathbb{R}.$
We view elements of $L$ as constant functions contained $\mathcal{O}(Y(x)_{L})$ and note that $(\pi_{L}|_{\psi_{\mathbb{C}}(x)})^{*}(L)=L\subseteq\mathbb{R}.$ For $i\in\{1,2\}$, we have $g_{i}|_{Y(x)_{L}}\in\mathcal{O}(Y(x)_{L})$. As
$(\pi_{L}|_{\psi_{\mathbb{C}}(x)})^{*}(g_{i}|_{Y(x)_{L}})=\pi_{L}^{*}(g_{i})(\psi_{\mathbb{C}}(x))\in L,$
the injectivity of $(\pi_{L}|_{\psi_{\mathbb{C}}(x)})^{*}$ implies that $g_{i}|_{Y(x)_{L}}\in L$, for $i=1,2$. So both $\pi_{L}^{*}(g_{1})$ and $\pi_{L}^{*}(g_{2})$ are constants on $Y(x)=\pi_{L}^{-1}(Y(x)_{L})\subseteq R(X)$, with values in $L$. We conclude that $h=g\circ\psi_{\mathbb{C}}=(\pi_{L}^{*}(g_{1})+\pi_{L}^{*}(g_{2})i)\circ\psi_{\mathbb{C}}$ is constant on $Z^{\mathbb{R}}(x)$, since $Z^{\mathbb{R}}(x)=\psi_{\mathbb{C}}^{-1}(Y(x)(\mathbb{R}))$. ∎
@@ 2.4. Moduli space of rational maps
For $d\geq 2$, let ${\rm Rat}_{d}$ be the space of degree $d$ endomorphisms on $\mathbb{P}^{1}$. It is a smooth irreducible affine variety of dimension $12d+$ 1 [25]. The group ${\rm PGL}_{2}={\rm Aut}(\mathbb{P}^{1})$ acts on ${\rm Rat}_{d}$ by conjugacy. The geometric quotient
@@
is the (coarse) moduli space of endomorphisms of degree $d$ on $\mathbb{P}^{1}$, which is an irreducible affine variety of dimension $22d-$ 2, see [25, 24] for a detailed treatment. Let
@@
be the quotient morphism. Let ${\rm FL}_{d}\subseteq{\rm Rat}_{d}$ be the locus of flexible Lattès maps, which is Zariski closed in ${\rm Rat}_{d}$. Set $[{\rm FL}_{d}]:=\Psi({\rm FL}_{d})$. The flexible-Lattès locus $[{\rm FL}_{d}]$ is non-empty if and only if $d$ is a square, and in this case $[{\rm FL}_{d}]$ is equidimensional of dimension $1$.
The above construction works over any algebraically closed field of characteristic 00 and commutes with base changes.
In the proof of Theorem 2.1, we need to consider only repelling periodic points of rational maps and the notion of moduli spaces with certain marked periodic points is necessary. We introduce such moduli spaces here. See [11, §8] for more information.
@@
Next we introduce the moduli space of endomorphisms of degree $d$ with a finite sequence of marked periodic points whose periods satisfy certain conditions. Let $s_{1},\dots,s_{n}\in\mathbb{Z}_{\geq 0}$ be finitely many integers such that $s_{1}\leq\dots\leq s_{n}$ and $s_{i}\leq d^{i!}+1$ for $1\leq i\leq n$. We construct the space ${\rm R_{d}}(s_{1},\dots,s_{n})$ of rational maps of degree $d$ with $s_{n}$ marked $n!$-periodic points, such that among these, $s_{n-1}$ are already $(n-1)!$-periodic, $\dots$, such that among these $s_{2}$ marked $2!$-periodic points, $s_{1}$ are already $1$-periodic (i.e., fixed), and we always count with multiplicity, as follows. An $m$-periodic point of $f$ just means a fixed point of $f^{m}$ and we do not require its exact period to be $m$, hence an $m!$-periodic point is automatically a $k!$-periodic point for $k\geq m$. Consider the fiber product $({\rm Rat}_{d}[n!])^{s_{n}}_{/{\rm Rat}_{d}}$ of $s_{n}$ copies of ${\rm Rat}_{d}[n!]$ over ${\rm Rat}_{d}$. For $i\neq j\in\{1,\dots,s_{n}\}$, let
@@
be the projection to the $i,j$ coordinates. The diagonal $\Delta\subseteq({\rm Rat}_{d}[n!])^{2}_{/{\rm Rat}_{d}}$ is an irreducible component of $({\rm Rat}_{d}[n!])^{2}_{/{\rm Rat}_{d}}$. Consider the open subset
@@
in which the $s_{n}$ marked points are pairwise distinct. Let $U^{\prime}$ be the subset of $U$ of points $(f,x_{1},\dots,x_{s_{n}})$ satisfying $f^{m!}(x_{i})=x_{i}$ for all integers $1\leq m\leq n$ and $1\leq i\leq s_{m}$. It means that in $U^{\prime}$ the first $s_{m}$ points in the whole sequence of $s_{n}$ marked $n!$-periodic points are indeed $m!$-periodic, for all integers $1\leq m\leq n$. The set $U^{\prime}$ is open and closed in $U$. We then define ${\rm R_{d}}(s_{1},\dots,s_{n})$ to be the Zariski closure of $U^{\prime}$ in $({\rm Rat}_{d}[n!])^{s_{n}}_{/{\rm Rat}_{d}}$. For all integers $0\leq m\leq n$, define
@@
Recall that for a rational map $g(z)\in\mathbb{C}(z)$ of degree $\geq 2$ and a fixed point $z_{0}\in\mathbb{P}^{1}(\mathbb{C})$ of $g$, the multiplicity of the fixed point $z_{0}$ of $g$ is the order of $g(z)-z=0$ at $z_{0}$ when $z_{0}\neq\infty$, and we can define the multiplicity using a conjugation to move $z_{0}$ to $\mathbb{C}$ when $z_{0}=\infty$. For every $k\geq 1$, it is clear that a fixed point $x$ of $f_{t}^{k}$ has multiplicity $1$ iff
@@
to be the moduli space of endomorphisms of degree $d$ on $\mathbb{P}^{1}$ with $s_{n}$ marked $n!$-periodic points, such that among these, $s_{n-1}$ are already $(n-1)!$-periodic, $\dots$, such that in these $s_{2}$ marked $2!$-periodic points, $s_{1}$ are already $1$-periodic (i.e., fixed), and we always count with multiplicity. The moduli space $\mathcal{M}_{d}(s_{1},\dots,s_{n})$ is affine, in fact, so the construction is relatively easy. The morphisms $\phi_{n,m}$ and $\lambda_{s_{1},\dots,s_{n}}$ descend to
$[\phi_{n,m}]:\mathcal{M}_{d}(s_{1},\dots,s_{n})\to\mathcal{M}_{d}(s_{1},\dots,s_{m})$
@@
$[\phi_{n,0}]:\mathcal{M}_{d}(s_{1},\dots,s_{n})\to\mathcal{M}_{d},$
@@
$[\lambda_{s_{1},\dots,s_{n}}]:\mathcal{M}_{d}(s_{1},\dots,s_{n})\to\mathbb{A}^{s_{n}}.$
Then $[\phi_{n,0}]$ is étale at every point $x\in[\lambda_{s_{1},\dots,s_{n}}]^{-1}((\mathbb{A}^{1}\setminus\{1\})^{s_{n}})$.
All the schemes ${\rm Rat}_{d}$, $\mathcal{M}_{d}$, ${\rm R_{d}}(s_{1},\dots,s_{n})$, and $\mathcal{M}_{d}(s_{1},\dots,s_{n})$ which we consider are defined over $\mathbb{Q}$.
@@ 2.5. Length maps
Recall that $K$ is an algebraically closed subfield of $\mathbb{C}$ which is invariant under the complex conjugation. We use the terminology and notations in §2.1-2.4. For $d\geq 2$, let $s_{1},\dots,s_{n}\in\mathbb{Z}_{\geq 0}$ be finitely many integers such that $s_{1}\leq\dots\leq s_{n}$ and $s_{i}\leq d^{i!}+1$ for $1\leq i\leq n$. Let
@@
$[\lambda_{s_{1},\dots,s_{n}}]:\mathcal{M}_{d}(s_{1},\dots,s_{n})(\mathbb{C})\to\mathbb{C}^{s_{n}}$
and the norm map
@@ 2.6. Rigidity of length spectrum
Let $f$ be an endomorphism of $\mathbb{P}^{1}(\mathbb{C})$ of degree $d\geq 2$. As in [12, §8.3] the length spectrum
@@
of $f$ is a sequence of finite multisets [^1] A multiset is a set except allowing multiple instances for each of its elements. The number of the instances of an element is called the multiplicity. For example: $\{a,a,b,c,c,c\}$ is a multiset of cardinality $6$, the multiplicities for $a,b,c$ are 2,1,3, respectively., where $L(f)_{n}:=L_{n}(f)$ is the multiset of norms of multipliers (i.e., lengths) of all fixed points of $f^{n}$. In particular, $L(f)_{n}$ is a multiset of non-negative real numbers of cardinality $d^{n}+1$. For every $n\geq 1$, let $RL(f)_{n}$ be the sub-multiset of $L(f)_{n}$ consisting of all elements $>1$. We call
@@
is increasing, since every $n!$-periodic point is $(n+1)!$-periodic and the multiplicity of a repelling fixed point of $f^{n!}$ is $1$ by (2.4). Also, the sequence $(d^{n!}+1-\#RL^{*}(f)_{n})_{n\geq 1}$ is bounded from above, since every rational map on $\mathbb{P}^{1}(\mathbb{C})$ of degree $d\geq 2$ has only finitely many periodic cycles whose multiplier has norm $\leq 1$ (see [19, Theorem 13.1]). As $L(f),RL(f)$, and $RL^{*}(f)$ are invariant under conjugacy, they descend on $\mathcal{M}_{d}(\mathbb{C})$. For every $[f]\in\mathcal{M}_{d}(\mathbb{C})$, define
@@
Let $\Omega$ be the set of sequences $\{A_{n},n\geq 1\}$ of multisets consisting of real numbers of norm strictly larger than $1$ satisfying #$\#A_{n}\leq d^{n!}+1$ for $n\geq 1$, and such that for every $a\in A_{n}$ with multiplicity $m\geq 1$, we have $a^{n+1}\in A_{n+1}$ with multiplicity $\geq m$. For $A,B\in\Omega$, we write $A\subseteq B$ if $A_{n}\subseteq B_{n}$ as multisets for every $n\geq 1$. An element $A=\{A_{n}\}\in\Omega$ is called big if $(d^{n!}+1-\#A_{n})_{n=1}^{\infty}$ is bounded. For every endomorphism $f$ of $\mathbb{P}^{1}(\mathbb{C})$ of degree $d$, we have $RL^{*}(f)\in\Omega$ and it is big.
For a big $A\in\Omega$, Ji-Xie proved the following rigidity result on length spectrum, which is a main ingredient of the proof of Theorem 2.1.
@@
Let $f:\mathbb{P}_{\mathbb{C}}^{1}\to\mathbb{P}_{\mathbb{C}}^{1}$ be a rational map of degree $d\geq 2$. Assume that $f$ is not a flexible Lattès map and for every $x\in{\rm Per}(f)(\mathbb{C})$, $|\rho_{f}(x)|\in K$. We want to show that $[f]\in\mathcal{M}_{d}(\mathbb{C})$ is not transcendental over $K$ for the model $(\mathcal{M}_{d})_{K}$. Now assume that $[f]$ is transcendental.
@@
$A_{n}=\{|\rho(x_{i})|^{n!}:i=1,\dots,s_{n}\}.$
Let $[f_{n}]\in\mathcal{M}_{d}(s_{1},\dots,s_{n})(\mathbb{C})$ be the point presented by $(f,x_{1},\dots,x_{s_{n}})$. It is clear that $[\phi_{n,0}]([f_{n}])=[f]$ for every $n\geq 1$.
Since $[f]$ is transcendental, for every $n\geq 1$, $[f_{n}]$ is transcendental as well. By Lemma 2.14, we get $\dim_{\mathbb{R}}(Z^{\mathbb{R}}([f_{n}]))\geq 1$ for every $n\geq 1$. Our assumption implies that
@@
Then from the last paragraph of §2.5, we see that $|\lambda_{s_{1},\dots,s_{n}}|$ is constant on $Z^{\mathbb{R}}(f_{n})$. As $|\lambda_{s_{1},\dots,s_{n}}|([f_{n}])\in(1,+\infty)^{s_{n}}$, the morphism $[\phi_{n,0}]$ is étale in a neighborhood of $Z^{\mathbb{R}}(f_{n})$. Since $[\phi_{n,0}]$ is a finite morphism, the subset
@@
Then for every $g\in V_{N}$, we have $A\subseteq RL^{*}([g])$. Since $[f]\not\in[{\rm FL}_{d}]$, the set $Z^{\mathbb{R}}([f_{N}])$ is real irreducible with $\dim_{\mathbb{R}}Z^{\mathbb{R}}([f_{N}])\geq 1$, hence $V_{N}\cap(\mathcal{M}_{d}(\mathbb{C})\setminus[{\rm FL}_{d}])$ is infinite. This contradicts to Theorem 2.16, which concludes the proof. ∎

### v3

@@ 2.3. Transcendental points
Write $h=g\circ\psi_{\mathbb{C}}$, where $g\in\mathcal{O}^{L}(R(X))\otimes_{L}K$. Write $g=\pi_{L}^{*}(g_{1})+\pi_{L}^{*}(g_{2})i$, where $g_{1},g_{2}\in\mathcal{O}(R(X_{K}))$. Since $h(x)\in K$, both $\pi_{L}^{*}(g_{1})(\psi_{\mathbb{C}}(x))$ and $\pi_{L}^{*}(g_{2})(\psi_{\mathbb{C}}(x))$ are in $L$. The map $\pi_{L}\mid_{\psi_{\mathbb{C}}(x)}:\psi_{\mathbb{C}}(x)\to Y(x)_{L}$ induces an injection
$(\pi_{L}\mid_{\psi_{\mathbb{C}}(x)})^{*}:\mathcal{O}(Y(x)_{L})\hookrightarrow\mathbb{R}.$
We view elements of $L$ as constant functions in $\mathcal{O}(Y(x)_{L})$ and note that
$(\pi_{L}\mid_{\psi_{\mathbb{C}}(x)})^{*}(L)=L\subseteq\mathbb{R}.$
For $i\in\{1,2\}$, we have $g_{i}\mid_{Y(x)_{L}}\in\mathcal{O}(Y(x)_{L})$. Since
$(\pi_{L}\mid_{\psi_{\mathbb{C}}(x)})^{*}(g_{i}\mid_{Y(x)_{L}})=\pi_{L}^{*}(g_{i})(\psi_{\mathbb{C}}(x))\in L,$
the injectivity of $(\pi_{L}\mid_{\psi_{\mathbb{C}}(x)})^{*}$ implies that $g_{i}\mid_{Y(x)_{L}}\in L$ for $i=1,2$. So both $\pi_{L}^{*}(g_{1})$ and $\pi_{L}^{*}(g_{2})$ are constant on $Y(x)=\pi_{L}^{-1}(Y(x)_{L})\subseteq R(X)$, with values in $L$. We conclude that $h=g\circ\psi_{\mathbb{C}}=(\pi_{L}^{*}(g_{1})+\pi_{L}^{*}(g_{2})i)\circ\psi_{\mathbb{C}}$ is constant on $Z^{\mathbb{R}}(x)$, since $Z^{\mathbb{R}}(x)=\psi_{\mathbb{C}}^{-1}(Y(x)(\mathbb{R}))$. ∎
@@ 2.4. Moduli space of rational maps
For $d\geq 2$, let ${\rm Rat}_{d}$ be the space of degree $d$ endomorphisms on $\mathbb{P}^{1}$. It is a smooth irreducible affine variety of dimension $2d+1$ [25]. The group ${\rm PGL}_{2}={\rm Aut}(\mathbb{P}^{1})$ acts on ${\rm Rat}_{d}$ by conjugacy. The geometric quotient
@@
is the (coarse) moduli space of endomorphisms of degree $d$ on $\mathbb{P}^{1}$, which is an irreducible affine variety of dimension 2d22d-2, see [25, 24] for a detailed treatment. Let
@@
be the quotient morphism. Let ${\rm FL}_{d}\subseteq{\rm Rat}_{d}$ be the locus of flexible Lattès maps, which is Zariski closed in ${\rm Rat}_{d}$. Set ${\rm FL}_{d}]:=\Psi({\rm FL}_{d})$. The flexible-Lattès locus ${\rm FL}_{d}]$ is non-empty if and only if $d$ is a square, and in this case ${\rm FL}_{d}]$ is equidimensional of dimension $1$.
The above construction works over any algebraically closed field of characteristic 0 and commutes with base changes.
In the proof of Theorem 2.1, we need to consider only repelling periodic points of rational maps, and the notion of moduli spaces with certain marked periodic points is necessary. We introduce such moduli spaces here. See [11, §8] for more details.
@@
Next, we introduce the moduli space of endomorphisms of degree $d$ with a finite sequence of marked periodic points whose periods satisfy certain conditions. Let $s_{1},\dots,s_{n}\in\mathbb{Z}_{\geq 0}$ be finitely many integers such that $s_{1}\leq\dots\leq s_{n}$ and $s_{i}\leq d^{i!}+1$ for $1\leq i\leq n$. We construct the space ${\rm R_{d}}(s_{1},\dots,s_{n})$ of rational maps of degree $d$ with $s_{n}$ marked $n!$-periodic points, such that among these, $s_{n-1}$ are already (n1)!(n-1)!-periodic, $\dots$, and among these $s_{2}$ marked $2!$-periodic points, $s_{1}$ are already $1$-periodic (i.e., fixed), always counted with multiplicity, as follows. An $m$-periodic point of $f$ means a fixed point of $f^{m}$, and we do not require its exact period to be $m$; hence, an $m!$-periodic point is automatically $k!$-periodic for $k\geq m$. Consider the fiber product $({\rm Rat}_{d}[n!])^{s_{n}}_{/{\rm Rat}_{d}}$ of $s_{n}$ copies of ${\rm Rat}_{d}[n!]$ over ${\rm Rat}_{d}$. For $i\neq j\in\{1,\dots,s_{n}\}$, let
@@
be the projection to the $i$-th and $j$-th coordinates. The diagonal $\Delta\subseteq({\rm Rat}_{d}[n!])^{2}_{/{\rm Rat}_{d}}$ is an irreducible component of $({\rm Rat}_{d}[n!])^{2}_{/{\rm Rat}_{d}}$. Consider the open subset
@@
in which the $s_{n}$ marked points are pairwise distinct. Let $U^{\prime}$ be the subset of $U$ consisting of points $(f,x_{1},\dots,x_{s_{n}})$ satisfying $f^{m!}(x_{i})=x_{i}$ for all integers $1\leq m\leq n$ and $1\leq i\leq s_{m}$. This means that in $U^{\prime}$, the first $s_{m}$ points in the whole sequence of $s_{n}$ marked $n!$-periodic points are indeed $m!$-periodic, for all integers $1\leq m\leq n$. The set $U^{\prime}$ is open and closed in $U$. We then define ${\rm R_{d}}(s_{1},\dots,s_{n})$ to be the Zariski closure of $U^{\prime}$ in $({\rm Rat}_{d}[n!])^{s_{n}}_{/{\rm Rat}_{d}}$. For all integers $0\leq m\leq n$, define
@@
Recall that for a rational map $g(z)\in\mathbb{C}(z)$ of degree $\geq 2$ and a fixed point $z_{0}\in\mathbb{P}^{1}(\mathbb{C})$ of $g$, the multiplicity of the fixed point $z_{0}$ of $g$ is the order of g(z)z=0g(z)-z=0 at $z_{0}$ when $z_{0}\neq\infty$, and we can define the multiplicity using a conjugation to move $z_{0}$ to $\mathbb{C}$ when $z_{0}=\infty$. For every $k\geq 1$, it is clear that a fixed point $x$ of $f_{t}^{k}$ has multiplicity $1$ if and only if
@@
to be the moduli space of endomorphisms of degree $d$ on $\mathbb{P}^{1}$ with $s_{n}$ marked $n!$-periodic points, such that among these, $s_{n-1}$ are already (n1)!(n-1)!-periodic, $\dots$, and among these $s_{2}$ marked $2!$-periodic points, $s_{1}$ are already $1$-periodic (i.e., fixed), always counted with multiplicity. The moduli space $\mathcal{M}_{d}(s_{1},\dots,s_{n})$ is affine, which makes the construction relatively easy. The morphisms $\phi_{n,m}$ and $\lambda_{s_{1},\dots,s_{n}}$ descend to
$\phi_{n,m}]:\mathcal{M}_{d}(s_{1},\dots,s_{n})\to\mathcal{M}_{d}(s_{1},\dots,s_{m})$
@@
$\phi_{n,0}]:\mathcal{M}_{d}(s_{1},\dots,s_{n})\to\mathcal{M}_{d},$
@@
$\lambda_{s_{1},\dots,s_{n}}]:\mathcal{M}_{d}(s_{1},\dots,s_{n})\to\mathbb{A}^{s_{n}}.$
Then $\phi_{n,0}]$ is étale at every point $x\in[\lambda_{s_{1},\dots,s_{n}}]^{-1}((\mathbb{A}^{1}\setminus\{1\})^{s_{n}})$.
All the schemes ${\rm Rat}_{d}$, $\mathcal{M}_{d}$, ${\rm R_{d}}(s_{1},\dots,s_{n})$, and $\mathcal{M}_{d}(s_{1},\dots,s_{n})$ considered here are defined over $\mathbb{Q}$.
@@ 2.5. Length maps
Recall that $K$ is an algebraically closed subfield of $\mathbb{C}$ which is invariant under complex conjugation. We use the terminology and notations from §2.12.4.
For $d\geq 2$, let $s_{1},\dots,s_{n}\in\mathbb{Z}_{\geq 0}$ be finitely many integers such that $s_{1}\leq\dots\leq s_{n}$ and $s_{i}\leq d^{i!}+1$ for $1\leq i\leq n$. Let
@@
$\lambda_{s_{1},\dots,s_{n}}]:\mathcal{M}_{d}(s_{1},\dots,s_{n})(\mathbb{C})\to\mathbb{C}^{s_{n}}$
and the absolute value map
@@ 2.6. Rigidity of length spectrum
Let $f$ be an endomorphism of $\mathbb{P}^{1}(\mathbb{C})$ of degree $d\geq 2$. As in [12, §8.3], the length spectrum
@@
of $f$ is a sequence of finite multisets [^1] A multiset is a set except allowing multiple instances for each of its elements. The number of instances of an element is called its multiplicity. For example: $\{a,a,b,c,c,c\}$ is a multiset of cardinality $6$, with multiplicities $2,1,3$ for $a,b,c$, respectively., where $L(f)_{n}:=L_{n}(f)$ is the multiset of lengths of all fixed points of $f^{n}$. In particular, $L(f)_{n}$ is a multiset of non-negative real numbers of cardinality $d^{n}+1$. For every $n\geq 1$, let $RL(f)_{n}$ be the sub-multiset of $L(f)_{n}$ consisting of all elements >1>1. We call
@@
is non-decreasing, since every $n!$-periodic point is $(n+1)!$-periodic and the multiplicity of a repelling fixed point of $f^{n!}$ is $1$ by (2.4). Also, the sequence $(d^{n!}+1-\#RL^{*}(f)_{n})_{n\geq 1}$ is bounded from above, since every rational map on $\mathbb{P}^{1}(\mathbb{C})$ of degree $d\geq 2$ has only finitely many periodic cycles whose multiplier has absolute value $\leq 1$ (see [18, Theorem 13.1]). Since $L(f),RL(f)$, and $RL^{*}(f)$ are invariant under conjugacy, they descend to $\mathcal{M}_{d}(\mathbb{C})$. For every $f]\in\mathcal{M}_{d}(\mathbb{C})$, define
@@
Let $\Omega$ be the set of sequences $\{A_{n},n\geq 1\}$ of multisets consisting of real numbers of absolute values >1>1 satisfying $A_{n}\leq d^{n!}+1$ for $n\geq 1$, and such that for every $a\in A_{n}$ with multiplicity $m\geq 1$, we have $a^{n+1}\in A_{n+1}$ with multiplicity $\geq m$. For $A,B\in\Omega$, we write $A\subseteq B$ if $A_{n}\subseteq B_{n}$ as multisets for every $n\geq 1$. An element $A=\{A_{n}\}\in\Omega$ is called big if $(d^{n!}+1-\#A_{n})_{n=1}^{\infty}$ is bounded. For every endomorphism $f$ of $\mathbb{P}^{1}(\mathbb{C})$ of degree $d$, we have $RL^{*}(f)\in\Omega$ and it is big.
For a big $A\in\Omega$, Ji-Xie proved the following rigidity result on length spectrum, which is a main ingredient in the proof of Theorem 2.1.
@@
Let $f:\mathbb{P}_{\mathbb{C}}^{1}\to\mathbb{P}_{\mathbb{C}}^{1}$ be a rational map of degree $d\geq 2$. Assume that $f$ is not a flexible Lattès map and that for every $x\in{\rm Per}(f)(\mathbb{C})$, $|\rho_{f}(x)|\in K$. We want to show that $f]\in\mathcal{M}_{d}(\mathbb{C})$ is not transcendental over $K$ for the model $(\mathcal{M}_{d})_{K}$. Now assume, for contradiction, that $[f]$ is transcendental.
@@
$A_{n}=\left\{\left|\rho_{f^{n!}}(x_{i})\right|:i=1,\dots,s_{n}\right\}.$
Let $f_{n}]\in\mathcal{M}_{d}(s_{1},\dots,s_{n})(\mathbb{C})$ be the point represented by $(f,x_{1},\dots,x_{s_{n}})$. It is clear that $\phi_{n,0}]([f_{n}])=[f]$ for every $n\geq 1$.
Since $[f]$ is transcendental, for every $n\geq 1$, $f_{n}]$ is also transcendental. By Lemma 2.14, we get $\dim_{\mathbb{R}}(Z^{\mathbb{R}}([f_{n}]))\geq 1$ for every $n\geq 1$. Our assumption implies that
@@
Then from the last paragraph of §2.5, we see that $|\lambda_{s_{1},\dots,s_{n}}|$ is constant on $Z^{\mathbb{R}}(f_{n})$. As $|\lambda_{s_{1},\dots,s_{n}}|([f_{n}])\in(1,+\infty)^{s_{n}}$, the morphism $\phi_{n,0}]$ is étale in a neighborhood of $Z^{\mathbb{R}}([f_{n}])$. Since $\phi_{n,0}]$ is a finite morphism, the subset
@@
Then for every $g\in V_{N}$, we have $A\subseteq RL^{*}([g])$. Since $f]\not\in[{\rm FL}_{d}]$, the set $Z^{\mathbb{R}}([f_{N}])$ is real irreducible with $\dim_{\mathbb{R}}Z^{\mathbb{R}}([f_{N}])\geq 1$, so $V_{N}\cap(\mathcal{M}_{d}(\mathbb{C})\setminus[{\rm FL}_{d}])$ is infinite. This contradicts Theorem 2.16, which completes the proof. ∎